# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 3.3 A positivity lemma
holds for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$.
@@
Here the term $O(1)$ is independent of $r$ and $n$. Combine (3.4.8), (3.4.10), (3.4), we have for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $r$),
@@
Here the term $O(1)$ depends only on $\varepsilon$. In particular, it is independent of $r$ and $n$. Finally combine (3.4.9) and (3.4), we get for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $r$),
@@
Here the term $O(1)$ is independent of $r,n$. For fixed $0<\varepsilon<\frac{c_{1}}{2c_{3}c_{4}}$, when $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ large enough, the term $O(1)$ in (3.4.13) is less than $\frac{1}{2}\left(c_{1}-2\varepsilon c_{3}c_{4}\right)T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$. Thus we have
@@
Remark 3.18.
Combined with Corollary 3.17, the inequality (3.4.14) can be intuitively understood as follows: the tangent points of $(s_{n},t_{n})(B)$ and ${\mathcal{D}}_{\rm red}$ account for a positive proportion of all intersection points, which has already been mentioned at the end of §1.2.
[…]
@@
For fixed $\varepsilon>0$, we can enlarge the set $\Sigma_{\varepsilon}$ of finite Lebesgue measure above such that for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$, the inequality
$N(r,\psi_{n},\widetilde{{\mathcal{R}}})\leq\varepsilon T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.4.15)
@@
Recall that $\widetilde{{\mathcal{R}}}=\widetilde{{\mathcal{R}}}_{0}+\widetilde{{\mathcal{R}}}_{1}$, where each irreducible component of $\widetilde{{\mathcal{R}}}_{0}$ dominates some component of $\widetilde{{\mathcal{R}}}$ and the image of $\widetilde{{\mathcal{R}}}_{1}$ in ${\mathcal{A}}\times_{\mathbb{C}}T$ is of codimension greater than 1. By definition, $N(r,\psi_{n},\widetilde{{\mathcal{R}}})=N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})+N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})$.
We first estimate $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})$. Since ${\mathcal{Z}}_{1}=(f\circ\pi)(\widetilde{{\mathcal{R}}}_{1})$, there is a positive integer $C_{1}$ such that $\widetilde{{\mathcal{R}}}_{1}\subseteq C_{1}(f\circ\pi)^{*}{\mathcal{Z}}_{1}$. Then we have
@@
Recall that we have a commutative diagram
@@
Since $\{\varphi_{n}\}_{n\geq 1}$ has a limit $\varphi_{\infty}$, $\{\varphi^{\prime}_{n}\}_{n\geq 1}$ also has a limit $\varphi^{\prime}_{\infty}$, which is a lift of $\varphi_{\infty}$ in ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$. Recall that $Z_{1}$ is the image of $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}_{\infty}}$ in ${\mathcal{A}}\times\{t_{\infty}\}$. By property (4) of the point $x$, we see that $\varphi_{\infty}(\mathbb{C})\nsubseteq Z_{1}$, and thus $\varphi^{\prime}_{\infty}(\mathbb{C})\nsubseteq{\mathcal{Z}}^{\prime}_{1}$. By Lemma 3.5, for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
@@
By our assumption, the Zariski closure of $\varphi_{\infty}(\mathbb{C})$ in ${\mathcal{A}}\times_{\mathbb{C}}T$ is $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$, whose intersection with $(Z_{1})_{b}$ has codimension greater than 1 in $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$ (see property (4) of $x$). By Lemma 3.6(2), we can enlarge $\Sigma_{\varepsilon}$ such that for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
$N(r,\varphi_{\infty},Z_{1})=N(r,\varphi_{\infty},(Z_{1})_{b})\leq\frac{\varepsilon}{4C_{1}C_{2}}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.4.20)

### v2

@@ 3.4 Step 2: lower bound of transversal intersections
holds for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$.
Upper bound of $N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})$
Now we establish the main lemma in this subsection, which gives an upper bound of $N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})$. This can also be understood as a lower bound for the transversal intersection numbers of $\varphi_{n}$ and ${\mathcal{D}}_{\rm red}$.
Lemma 3.16.
For fixed $\varepsilon>0$, we can enlarge the set $\Sigma_{\varepsilon}$ of finite Lebesgue measure above such that for any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$, the inequality
$N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})\leq\varepsilon T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.4.7)
holds for $n$ large enough ((depending on $\varepsilon$ and $r)$.
The key to the proof of Lemma 3.16 is the inequality (3.4.5) together with two limiting arguments: one is Lemma 3.7, and the other adapts the argument of Lemma 3.5. The philosophy is that, if the limit entire curve $\varphi_{\infty}$ has many transversal intersections with $({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$, then each $\varphi_{n}$ also has many transversal intersections with ${\mathcal{D}}_{\rm red}$.
The first limiting argument is contained in the following lemma. Recall that ${\mathcal{D}}_{t_{\infty}}^{\circ}\subseteq{\mathcal{D}}_{t_{\infty}}$ is a Zariski dense open subset of ${\mathcal{D}}_{t_{\infty}}$ satisfying the conditions in Lemma 3.7 (applying to ${\mathcal{D}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$ and the sequence $\{t_{n}\}_{n\geq 1}$). By property (3) of the set $S_{0}\subseteq B(\mathbb{C})$, we see that ${\mathcal{D}}_{b,t_{\infty}}^{\circ}$ is a Zariski dense open subset of ${\mathcal{D}}_{b,t_{\infty}}$.
Lemma 3.17.
Proof.
The assumption ${\rm ord}_{w_{n}}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}>1$ means that $({\rm d}\varphi^{\prime}_{n})_{0}(v_{\rm st})\in T_{\varphi^{\prime}_{n}(0)}{\mathcal{D}}_{\rm red}$ for all $n$ (large enough). Here $v_{\rm st}=\frac{\rm d}{{\rm d}z}$ is the tangent vector of ${\mathbb{D}}_{r^{\prime}_{n}}$ at 0 under the standard coordinate $z$. By the uniform converging property, $({\rm d}\varphi^{\prime}_{n})_{0}(v_{\rm st})$ tends to $({\rm d}\varphi^{\prime}_{\infty})_{0}(v_{\rm st})$. Since $y\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$, by Lemma 3.7, we see that $({\rm d}\varphi^{\prime}_{\infty})_{0}(v_{\rm st})\in T_{\varphi^{\prime}_{\infty}(0)}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Or equivalently, $({\rm d}\varphi_{\infty})_{z}(v_{\rm st})\in T_{\varphi_{\infty}(z)}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$, which contradicts to the assumption that ${\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$.
Now take a small neighbourhood $V_{z}$ of $z$ contained in ${\mathbb{D}}_{2|z|}$ such that $\varphi_{\infty}(V_{z})\subseteq U_{y}$. Use the uniformly converge again, we see that for $n$ large enough, $\varphi_{n}(V_{z})\subseteq U_{y}$. Therefore by the argument above, ${\rm ord}_{z^{\prime}}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}\leq 1$ holds for all $z^{\prime}\in V_{z}$ and sufficiently large $n$. This finishes the proof of the lemma. ∎
Now we are ready to prove Lemma 3.16. The argument is similar to that of Lemma 3.5, with slightly more elaborate estimates.
Proof of Lemma 3.16.
For fixed $r>0$ such that $\varphi_{\infty}(\partial{\mathbb{D}}_{r})\cap{\mathcal{D}}=\emptyset$ (we can enlarge $\Sigma_{\varepsilon}$ so that each $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$ satisfies this property), choose an open neighbourhood $W_{z}$ of $z$ (depending on $r$) for each $z\in\overline{{\mathbb{D}}}_{r}$ satisfying the following properties:
  1. (1)
  2. (2)
    $\log r-\log|z^{\prime}|\leq 2(\log r-\log|z|)$
  3. (3)
    If $|z|<1$, then $W_{z}\subseteq{\mathbb{D}}_{1}$.
  4. (4)
    If $\varphi_{\infty}(z)\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$ and ${\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$, then $W_{z}\subseteq V_{z}$, where $V_{z}$ is in Lemma 3.17.
By definition, we have
$\min\{\log r,\log r-\log\left|z\right|\}$ (3.4.8)
For $1\leq j\leq m$, define
By Rouché’s theorem again, for $n$ large enough, we have
Now for $n$ large enough and every $1\leq j\leq m$, if $\varphi_{\infty}(z_{j})\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$ and ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$, we see from Lemma 3.17 and assumption (4) that ${\rm ord}_{z}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}-\min\{1,{\rm ord}_{z}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}\}=0$ for all $z\in W_{z_{j}}$. Thus we have $S_{j}=0$. If ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=0$, then ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}=0$ so by the previous paragraph we also get $S_{j}=0$ for $n$ large enough.
for $n$ large enough.
Choose a positive integer $C>0$ such that $({\mathcal{D}}_{\rm red})_{b,t_{\infty}}\subseteq C({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Then we have
${\rm ord}_{z}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}\leq C{\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Moreover, if $z=z_{j}$ for some $j\in J_{2}$, then
${\rm ord}_{z}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}\leq C{\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=C\leq C{\rm ord}_{z}\varphi_{\infty}^{*}Z$. Hence by (3.4) we get
$\sum\limits_{j\in J_{1}}S_{j}+\sum\limits_{j\in J_{2}}S_{j}$ $\leq 2C\sum\limits_{j\in J_{1}}({\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\min\{\log r,\log r-\log|z_{j}|\}$
$+2C\sum\limits_{j\in J_{2}}({\rm ord}_{z_{j}}\varphi_{\infty}^{*}Z)\min\{\log r,\log r-\log|z_{j}|\}$
$\leq 2C\left(2(N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}))+N(r,\varphi_{\infty},Z)\right)$ (3.4.12)
By (3.4.5), with $\varepsilon$ replaced by $\frac{\varepsilon}{48C}$, we get
$N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq\frac{\varepsilon}{16C}N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
holds for all $r>0$ outside a set of finite Lebesgue measure. Enlarge $\Sigma_{\varepsilon}$ if necessary, we may assume that this holds for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$.
By Theorem 3.4 we easily see that $N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq 2T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ for $r$ sufficiently large. Thus we can enlarge $\Sigma_{\varepsilon}$ so that this holds for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$. Therefore we get
$N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq\frac{\varepsilon}{8C}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$ (3.4.13)
Denote $W=Z_{b}\cap(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. Then by our choice of $x$ (see property (3)), $W$ has codimension at least 2 in $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. By Theorem 3.6(2), we may enlarge $\Sigma_{\varepsilon}$ so that for any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,
$N(r,\varphi_{\infty},W)$ $\leq\frac{\varepsilon}{4C}T(r,\varphi_{\infty},(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})\cap({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
$=\frac{\varepsilon}{4C}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
Thus we have $N(r,\varphi_{\infty},Z)=N(r,\varphi_{\infty},W)\leq\frac{\varepsilon}{4C}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$. Combine this with (3.4.9), (3.4.11), (3.4), (3.4.13), we get
$N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})\leq\varepsilon T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
for any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $r$). This finishes the proof of Lemma 3.16. ∎
3.5 Step 3: upper bound of transversal intersections
On the other hand, we can estimate a lower bound for $N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})$ (or equivalently, an upper bound for the transversal intersection numbers of $\varphi_{n}$ with ${\mathcal{D}}_{\rm red}$). The precise estimate proceeds as follows.
Lemma 3.18.
There is a constant $c>0$ and a set $\Sigma\subseteq\mathbb{R}_{>0}$ of finite Lebesgue measure such that for all $r\in\mathbb{R}_{>0}\backslash\Sigma$, the inequality
$N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})\geq c\cdot T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.5.1)
holds for $n$ large enough ((depending on $r)$.
Proposition 3.1 follows immediately from Lemma 3.16 and Lemma 3.18. In fact, take $0<\varepsilon<\min\{\frac{1}{2},c\}$. Let $\Sigma_{\varepsilon}$ be the set of finite Lebesgue measure in Lemma 3.16. Take $r\in\mathbb{R}_{>0}\backslash(\Sigma_{\varepsilon}\cup\Sigma)$ large enough such that $T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})>0$. Then for $n$ large enough (depending on $\varepsilon$ and $r$) we get a contradiction from (3.4.7) and (3.5.1). Therefore, it remains to prove Lemma 3.18.
[…]
@@ 3.5 Step 3: upper bound of transversal intersections
Here the term $O(1)$ is independent of $r$ and $n$. Now for any $0<\varepsilon<\frac{1}{2}$, combine (3.4.6), (3.5.3), (3.5), we have for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $\varepsilon$ and $r$),
@@
Here the term $O(1)$ depends only on $\varepsilon$. In particular, it is independent of $r$ and $n$. Finally combine (3.5.2) and (3.5), we get for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $\varepsilon$ and $r$),
@@
Here the term $O(1)$ is independent of $r,n$. For fixed $0<\varepsilon<\frac{c_{1}}{2c_{3}c_{4}}$, when $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$ large enough, the term $O(1)$ in (3.5.6) is less than $\frac{1}{2}\left(\frac{c_{1}}{2}-2\varepsilon c_{3}c_{4}\right)T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$. Thus we have
@@
Tangency and a upper bound of $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$
Next we establish a relation between $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$ and $N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})$, the main result is Corollary 3.21. As we have mentioned in §1.3, an important point in this step is a tangency observation. Now we give the rigorous statement.
Recall that $\widetilde{{\mathcal{R}}}_{0}\subseteq\widetilde{{\mathcal{R}}}$ is the union of all irreducible components of $\widetilde{{\mathcal{R}}}$ which dominates an irreducible component of $\mathcal{R}$}.
Proposition 3.19.
Proof.
Corollary 3.20.
Proof.
We can translate Corollary 3.20 into an inequality in Nevanlinna theory. Recall that $\psi_{n},\varphi_{n}$ are respectively re-parametrizations of $\tilde{x}_{n},(s_{n},t_{n})$ defined in the last subsection. We have the following corollary.
Corollary 3.21.
There is a constant $c_{5}>0$ such that for all $n\geq 1$, the inequality
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})\leq c_{5}(N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red}))$ (3.5.8)
holds for all $r\geq 1$.
Proof.
Choose a positive integer $c_{0}>0$ such that $\widetilde{{\mathcal{R}}}_{0}\subseteq c(\widetilde{{\mathcal{R}}}_{0})_{\rm red}$. Then we have
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})\leq c_{0}N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})$ (3.5.9)
Corollary 3.20 just means that
In fact, if $\psi_{n}(z)\notin\widetilde{{\mathcal{R}}}_{0}$, then
${\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}=0\leq{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})-\min\{1,{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})\},$
If $\psi_{n}(z)\in\widetilde{{\mathcal{R}}}_{0}$, then
${\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}\leq{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})\quad{\rm and}\quad\min\{1,{\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}\}=\min\{1,{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})\}=1.$
So we always have
Hence by the definition of the counting function, we get (3.5.11).
Combine (3.5.9), (3.5.10) and (3.5.11) we see that
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$ $\leq c_{0}N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})$
$=c_{0}\left(N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})-N^{(1)}(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})+N^{(1)}(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})\right)$
$\leq 2c_{0}(N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})).$
Now (3.5.8) is true by taking $c_{5}=2c_{0}>0$. ∎
Upper bound of $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})$
Recall that $\widetilde{{\mathcal{R}}}=\widetilde{{\mathcal{R}}}_{0}+\widetilde{{\mathcal{R}}}_{1}$, and thus $N(r,\psi_{n},\widetilde{{\mathcal{R}}})=N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})+N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})$. We have established the estimates (3.5.7), (3.5.8) for $N(r,\psi_{n},\widetilde{{\mathcal{R}}}),N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$ respectively. Now we estimate the remaining term $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})$, which turns out to be negligible, due to the fact that the image of $\widetilde{{\mathcal{R}}}_{1}$ in $\mathcal{X}$} has codimension greater than 1.
@@
For fixed $\varepsilon>0$, we can enlarge the set $\Sigma_{\varepsilon}$ above such that for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$, the inequality
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})\leq\varepsilon T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.5.12)
@@
Since ${\mathcal{Z}}_{1}=(f\circ\pi)(\widetilde{{\mathcal{R}}}_{1})$, there is a positive integer $C_{1}$ such that $\widetilde{{\mathcal{R}}}_{1}\subseteq C_{1}(f\circ\pi)^{*}{\mathcal{Z}}_{1}$. Then we have
@@
Recall from “Choice of all data” that we have a commutative diagram
@@
Since $\{\varphi_{n}\}_{n\geq 1}$ has a limit $\varphi_{\infty}$, $\{\varphi^{\prime}_{n}\}_{n\geq 1}$ also has a limit $\varphi^{\prime}_{\infty}$, which is a lift of $\varphi_{\infty}$ in ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$. Recall that $Z_{1}$ is the image of $({\mathcal{Z}}^{\prime}_{1})_{t^{\prime}_{\infty}}$ in ${\mathcal{A}}\times\{t_{\infty}\}$. By property (4) of the point $x$, we see that $\varphi_{\infty}(\mathbb{C})\nsubseteq Z_{1}$, and thus $\varphi^{\prime}_{\infty}(\mathbb{C})\nsubseteq{\mathcal{Z}}^{\prime}_{1}$. By Lemma 3.5, for any $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,
@@
By our assumption, the Zariski closure of $\varphi_{\infty}(\mathbb{C})$ in ${\mathcal{A}}\times_{\mathbb{C}}T$ is $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$, whose intersection with $(Z_{1})_{b}$ has codimension greater than 1 in $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$ (see property (4) of $x$). By Lemma 3.6(2), we can enlarge $\Sigma_{\varepsilon}$ such that for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$,
$N(r,\varphi_{\infty},Z_{1})=N(r,\varphi_{\infty},(Z_{1})_{b})\leq\frac{\varepsilon}{2C_{1}C_{2}}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ (3.5.17)